Circular No. 8602

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005eb and 2005ec Further to IAUC 8598, N. Ponticello and W. Li report the LOSS/KAIT

discovery of an apparent supernova (designated 2005eb) on unfiltered KAIT images taken on Sept. 7.33 (at mag 18.4) and 16.33 UT (mag 18.8) and located at α = 0^{h}54^{m}50^{s}.94, δ = +29^{o}14′46′′.8 (equinox 2000.0), which is 9′′.0 east and 2′′.4 south of the center of UGC 556. Nothing was visible at this location on a KAIT image taken on Aug. 27.32 (limiting mag 19.5).

Ponticello and Li also report the LOSS/KAIT discovery of another supernova (designated SN 2005ec) on unfiltered KAIT images taken on Sept. 16.54 (at mag 16.8) and 17.54 UT (mag 16.6) and located at α = 4^{h}54^{m}19^{s}.17, δ = +1^{o}38′27′′.6 (equinox 2000.0), which is 1′′.3 west and 2′′.5 north of the center of NGC 1690. Nothing was visible at this location on a KAIT image taken on Jan. 11.13 (limiting mag 19.0).

The “Nearby Supernova Factory” group (G. Aldering et al.; cf. CBET 230) reports that a spectrum (range 320–1000 nm) of SN 2005ec, obtained on Sept. 21.6 UT with the Supernova Integral Field Spectrograph on the University of Hawaii 2.2-m telescope, shows it to be a type-Ia supernova at an age of two weeks after maximum and at a redshift of z = 0.03, in agreement with the host redshift (NGC 1690; Ostriker et al. 1988, A.J. 96, 1775). The spectrum is similar to that of SN 1994D at this phase, shifted to the host’s redshift.

V838 MONOCEROTIS M. Claussen, National Radio Astronomy Observatory (NRAO); K.

Healy and S. Starrfield, Arizona State University; and H. E. Bond, Space Telescope Science Institute, report SiO maser observations of the unusual eruptive variable V838 Mon using the NRAO Very Large Array on Sept. 1: “SiO maser emission was first detected by Deguchi et al. (2005, http://

xxx.lanl.gov/abs/astro-ph/?0507363) on 2005 Feb. 23, strengthening

by Apr. 24. Relative to the April observations, we find that the peak flux in the v=2, J=1-0 SiO transition at 42.82 GHz has increased further by a factor of four to 4.98 ± 0.01 Jy. The peak flux in the v=1, J=1-0 transition at 43.12 GHz has also strengthened by a factor of three to 2.83 ± 0.01 Jy. The maser emission in both transitions is unresolved (angular size < 0′′.3). Further monitoring with the VLA and the Very Long Baseline Array is planned, and monitoring is encouraged at other wavelengths, especially in the near- and mid-infrared.”

2005 September 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green